In 2019, Barry Grey Inc. sold 42,000 units at a selling price of $44 per unit. The company manufactured 70,000 units. Variable manufacturing costs were $22 per unit manufactured. Fixed manufacturing costs amounted to $334,000. Variable marketing costs were $9 per unit sold, and the budgeted and actual fixed marketing costs were $30,000. Other fixed operating expenses amounted to $22,000. There was no beginning inventory.

Round all answers to the nearest whole number.

a) Calculate the company's 2019 operating income using absorption costing.

Revenues      
Cost of Goods Sold:
Beginning Inventory
Cost of Goods Manufactured
Less: Ending Inventory
Cost of Goods Sold
Gross Margin
Operating Expenses:
Marketing Costs
Other Fixed Operating Expenses
Income from Operations


b) Calculate the company's 2019 operating income using variable costing.

Revenues      
Variable Cost of Goods Sold
Beginning Inventory
Cost of Goods Manufactured
Less: Ending Inventory
Variable Cost of Goods Sold
Variable Marketing Costs
Contributed Margin
Operating Expenses:
Fixed Manufacturing Overhead Costs
Fixed Marketing Costs
Other Fixed Operating Expenses
Income from Operations
Answer #1

Answer

  • Requirement 1

Revenues

  

$1,848,000

Cost of Goods Sold:

Beginning Inventory

$0

Cost of Goods Manufactured

$1,874,000

Less: Ending Inventory

$749,600

Cost of Goods Sold

$1,124,400

Gross Margin

$723,600

Operating Expenses:

Marketing Costs

$408,000

Other Fixed Operating Expenses

$22,000

$430,000

Income from Operations

$293,600

--Workings

Revenues

  

=42000*44

Cost of Goods Sold:

Beginning Inventory

0

Cost of Goods Manufactured

=(70000*22)+334000

Less: Ending Inventory

=1874000-1124400

Cost of Goods Sold

=+(42000*22)+(42000*334000/70000)

Gross Margin

=1848000-1124400

Operating Expenses:

Marketing Costs

=+(42000*9)+30000

Other Fixed Operating Expenses

22000

=408000+22000

Income from Operations

=723600-430000

  • Requirement 2: Variable costing

.

Revenues

  

$1,848,000

Variable Cost of Goods Sold

Beginning Inventory

$0

Cost of Goods Manufactured

$1,540,000

Less: Ending Inventory

$616,000

Variable Cost of Goods Sold

$924,000

Variable Marketing Costs

$378,000

$1,302,000

Contributed Margin

$546,000

Operating Expenses:

Fixed Manufacturing Overhead Costs

$334,000

Fixed Marketing Costs

$30,000

Other Fixed Operating Expenses

$22,000

$386,000

Income from Operations

$160,000

--Working

.

Revenues

  

=42000*44

Variable Cost of Goods Sold

Beginning Inventory

0

Cost of Goods Manufactured

=70000*22

Less: Ending Inventory

=1540000-924000

Variable Cost of Goods Sold

=42000*22

Variable Marketing Costs

=42000*9

=924000+378000

Contributed Margin

=1848000-1302000

Operating Expenses:

Fixed Manufacturing Overhead Costs

334000

Fixed Marketing Costs

30000

Other Fixed Operating Expenses

22000

=334000+30000+22000

Income from Operations

=546000-386000